A parabolic reflector is used by a sports broadcast engineer to amplify the conversations between football players on the field. If the reflector has the shape x = 5y2, where should the microphone be placed?

Answer by lwsshak3(11628)

A parabolic reflector is used by a sports broadcast engineer to amplify the conversations between football players on the field. If the reflector has the shape x = 5y2, where should the microphone be placed?
***
x=5y^2
y^2=(1/5)x
this is an equation of a parabola that opens rightward with vertex at the origin
Its basic equation: y^2=4px
For given equation:
vertex:(0,0)
axis of symmetry: y=0
4p=1/5
p=1/20
focal point:(1/20,0) (p-distance from vertex on the axis of symmetry)
where should the microphone be placed? at the focal point (1/20, 0)